# Just Ramp-up: Unleash the Potential of Regression-based Estimator for A/B Tests under Network Interference

### v1

@@ A.3 Proof of Theorem 2
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-\frac{d_{1}+d_{2}}{2n}\mathbf{1})^{T}\right]\begin{pmatrix}B&0\\ 0&B\end{pmatrix}\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\operatorname{trace}\left(\left(\mathbb{E}\left[\mathbf{z}\mathbf{z^{T}}\right]-\frac{d_{1}+d_{2}}{2n}\mathbb{E}\left[\mathbf{z}\mathbf{1}^{T}\right]\right)\begin{pmatrix}B&0\\ 0&B\end{pmatrix}\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\left(\operatorname{trace}\left((\mathbb{E}[\mathbf{z}_{1}\mathbf{z}_{1}^{T}]+\mathbb{E}[\mathbf{z}_{2}\mathbf{z}_{2}^{T}])B\right)-\frac{d_{1}+d_{2}}{2n}(\frac{d_{1}}{n}+\frac{d_{2}}{n})\operatorname{trace}\left(\mathbf{1}\mathbf{1}^{T}B\right)\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\left(\frac{d_{1}(d_{1}-1)+d_{2}(d_{2}-1)}{n(n-1)}-\frac{(d_{1}+d_{2})^{2}}{2n^{2}}\right)\operatorname{trace}\left(\mathbf{1}\mathbf{1}^{T}B\right)$
@@ A.6 Proof of Lemma 1
For the second case, we will prove that the two coefficients of term $(\sum_{ij}B_{ij})^{2}$ in the expression of $\operatorname{Var}((\mathbf{z}-w\mathbf{1})^{T}B\mathbf{z})$ can cancel out. Namely, we will prove:
$\mathbb{E}[(z_{i}-w)(z_{k}-w)z_{j}z_{l}]-\left(\mathbb{E}[(\mathbf{z}-w\mathbf{1})^{T}B\mathbf{z}]/\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)\right)^{2}=O(\frac{1}{n})$ (97)
@@
$\frac{\mathbb{E}[(\mathbf{z}-w\mathbf{1})^{T}B\mathbf{z}]}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)}$ $=\frac{\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-w\mathbf{1})^{T}\right]B\right)}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)}$ (100)
$=\frac{\operatorname{trace}\left((\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]-cw\mathbf{1}\mathbf{1}^{T})B\right)}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)}$
@@
In the last equation, we again utilize $B_{ii}=0$ for all unit $i\in[n]$. Combine two equations above, we successfully control the order of coefficient of term $(\sum_{ij}B_{ij})^{2}$, in the expression of concerned variance $\operatorname{Var}((\mathbf{z}-w\mathbf{1})^{T}B\mathbf{z})$.
@@ A.7 Proof of Theorem 4
$\operatorname{Var}\left(\sum_{t=1}^{T}\left(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1}\right)^{T}B\mathbf{z}_{t}\right)=\sum_{t=1}^{T}\operatorname{Var}\left(\left(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1}\right)^{T}B\mathbf{z}_{t}\right)=O(\frac{1}{n})$ (101)
@@ A.8 Proof of Theorem 5
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ (103)
$+2\sum_{t_{1}<t_{2}}\operatorname{Cov}$ $[(\mathbf{z}_{t_{1}}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t_{1}},(\mathbf{z}_{t_{2}}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t_{2}}]$
@@ B.1 Parameters
Moreover, we implement Monte Carlo simulations with 1,000 continuous random seeds, which control the treatment allocation. We fix another random seed that controls the initialization of the graph neural network, among all cases examined in our paper. If such a random seed also changes in a Monte Carlo simulation, empirically, there would be about ten percent of instances where the model falls into a local minimum in the very early stage.
@@ B.2 Performance of Linear Regression Estimator
We observe that the bias reductions introduced by merging compared to the case of $t=1$ verify

### v2

@@ A.3 Proof of Theorem 2
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-\frac{d_{1}+d_{2}}{2n}\mathbf{1})^{\top}\right]\begin{pmatrix}B&0\\ 0&B\end{pmatrix}\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\operatorname{trace}\left(\left(\mathbb{E}\left[\mathbf{z}\mathbf{z^{\top}}\right]-\frac{d_{1}+d_{2}}{2n}\mathbb{E}\left[\mathbf{z}\mathbf{1}^{\top}\right]\right)\begin{pmatrix}B&0\\ 0&B\end{pmatrix}\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\left(\operatorname{trace}\left((\mathbb{E}[\mathbf{z}_{1}\mathbf{z}_{1}^{\top}]+\mathbb{E}[\mathbf{z}_{2}\mathbf{z}_{2}^{\top}])B\right)-\frac{d_{1}+d_{2}}{2n}(\frac{d_{1}}{n}+\frac{d_{2}}{n})\operatorname{trace}\left(\mathbf{1}\mathbf{1}^{\top}B\right)\right)$
$=\frac{2n}{(d_{1}+d_{2})(2n-(d_{1}+d_{2}))}\left(\frac{d_{1}(d_{1}-1)+d_{2}(d_{2}-1)}{n(n-1)}-\frac{(d_{1}+d_{2})^{2}}{2n^{2}}\right)\operatorname{trace}\left(\mathbf{1}\mathbf{1}^{\top}B\right)$
@@ A.6 Proof of Lemma 1
For the second case, we will prove that the two coefficients of term $(\sum_{ij}B_{ij})^{2}$ in the expression of $\operatorname{Var}((\mathbf{z}-w\mathbf{1})^{\top}B\mathbf{z})$ can cancel out. Namely, we will prove:
$\mathbb{E}[(z_{i}-w)(z_{k}-w)z_{j}z_{l}]-\left(\mathbb{E}[(\mathbf{z}-w\mathbf{1})^{\top}B\mathbf{z}]/\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)\right)^{2}=O(\frac{1}{n})$ (97)
@@
$\frac{\mathbb{E}[(\mathbf{z}-w\mathbf{1})^{\top}B\mathbf{z}]}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)}$ $=\frac{\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-w\mathbf{1})^{\top}\right]B\right)}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)}$ (100)
$=\frac{\operatorname{trace}\left((\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]-cw\mathbf{1}\mathbf{1}^{\top})B\right)}{\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)}$
@@
In the last equation, we again utilize $B_{ii}=0$ for all unit $i\in[n]$. Combine two equations above, we successfully control the order of coefficient of term $(\sum_{ij}B_{ij})^{2}$, in the expression of concerned variance $\operatorname{Var}((\mathbf{z}-w\mathbf{1})^{\top}B\mathbf{z})$.
@@ A.7 Proof of Theorem 4
$\operatorname{Var}\left(\sum_{t=1}^{T}\left(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1}\right)^{\top}B\mathbf{z}_{t}\right)=\sum_{t=1}^{T}\operatorname{Var}\left(\left(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1}\right)^{\top}B\mathbf{z}_{t}\right)=O(\frac{1}{n})$ (101)
@@ A.8 Proof of Theorem 5
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ (103)
$+2\sum_{t_{1}<t_{2}}\operatorname{Cov}$ $[(\mathbf{z}_{t_{1}}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t_{1}},(\mathbf{z}_{t_{2}}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t_{2}}]$
@@ B.1 Parameters
Moreover, we conduct Monte Carlo simulations with 1,000 continuous random seeds to control treatment allocation and randomness of potential outcomes. For all cases examined in this paper, we fix another random seed (set to 2) to control the random initialization of the graph neural network. Empirically, if this random seed were allowed to vary during the Monte Carlo simulations, approximately ten percent of instances would fall into a local minimum in the early stages of training.
@@ B.2 Performance of Linear Regression Estimator
We observe that the bias reductions introduced by merging compared to the case of $t=1$ verify our theoretical results.